Logbooks general detector daq Analysis Physics 12 GeV HallA OldDVCS2 DVCS2Ana Parity
runliste issues procurement survey minutes web compton2 NewIssues NewProcurement
  Minutes  Not logged in ELOG logo
Message ID: 11     Entry time: 07/30/04, 09:42
Author: Franck Sabatie 
Subject: DVCS meeting minutes - 07/28/04 
  Present: A. Camsonne, J-M. Fieschi, A. Chanteleauze, F. Sabatie, C.
Munoz Camacho, M. Beaumel, C.E. Hyde-Wright, D. Hayes, G. Gavalian, S.
Nanda, K. McKormick, J. Segal, E. Voutier, M. Mazouz

  -> Jack and everyone/Installation:

  - 08/02 :         DAQ side un-cabling. Preparation for transport.

  - 08/03 :         Move DVCS electronics to Hall A.

  - 08/03-09 :    * Plug-in DVCS DAQ to Hall A DAQ with Bob's help.
                  * Detector side un-cabling in clean room.

  - 08/09 :         Move DVCS detectors in Hall A.

  - 08/10-09/03 : * Unstack calorimeter, repair, stack again.
                  * Standalone DAQ tests.
                  * Pull RG213 from spectrometer hutt down to pivot.
                  * Slow control + HV cabling (decide where to put it).
                  * Patch Panels installation. Plug in RG213 and DAQ side.
                  * Write security documents.
                  * DVCS detector on platform on 09/03.

  - 09/04-09/16 : * Plug-in cables on the detector (RG213, etc).
                  * DAQ timing with Bob Michael.
                  * Detectors check out (about 8-10 days of tests w/ DAQ).

  Also discussed a few minor issues with Jack.

  -> Charles and Gagik/Proton-Array:

  Charles presented a possibility to actually look for elastic protons
into the proton array while doing the spectrometer check-out with HRSleft
(e-) at 20 degrees. To illuminate two towers, one can do different
settings moving back the DVCS stack by a total of about 80cm at the
furthest setting. We all agreed it would be beneficial to do such a quick
test while doing the HRSleft check-out.

  Gagik showed latest results on the P-A. Everything looks okay except one
block (badly located) which appears to be dead (most likely PMT). This
will be investigated when the DVCS detector is in the Hall.

  -> Eric/VETO:

  The HV short was investigated to come from patch panel. Electronics
group will help repair the electronic boards of the patch panel. Should be
done before end of next week.

  Other than that, everything looks fine.

  -> Carlos/Calorimeter

  Showed new results about block positioning (see ELog for details).
Determination of this block center is accurate to within 1mm2, but after
investigation, it seems that the amplitude dependency versus a move of
around 1mm is 5% (which remains to be understood). The LED method seems to
have a systematic error at least that size just because of positioning
error. With new position from scans (vastly improved the X-Y matrix of
block position), the discrepancy between last cosmics gain-matching and a
new LED run goes down to about 9% which is the record so far. A vote of
what method would be closer to the real electron or photon events during
the experiment ended up largely in favor of the LED method, we will see :)
Work on the broken blocks (broken solderings actually) will start as soon
as tuesday next week. Kathy has agreed to setup a table to work on this in
the clean room.

  -> Kathy/Slow control

  Kathy figured out the right setup for the VME scan ADCs and tested them.
Everything looks good. P-A GUI is there, Calorimeter GUI is there, just
miss the Veto GUI. The Veto guyz will interact with Kathy so that the GUI
is done. Good news for Kathy and bad news for us, Kathy got a job in
Washington State! Congratulations Kathy !!!

  -> Sirish/TSOP and misc items

  Standard operating procedures need to be written shortly. Sirish will
start and spokespersons will fill in the blanks.

  A survey meeting is absolutely necessary. We will try to schedule it as
soon as Pierre is back (August 5th).

  Some money remains to be spent, if you have items, let him know...
ELOG V3.1.4-unknown